# Complexity at Scale: A Quantitative Analysis of an Alibaba Microservice Deployment

### v2

@@ Abstract.
Recent studies have begun to explore the characteristics of real-world large-scale microservice deployments. However, their operational complexities, and the degree to which these complexities are consistent across different deployments, remains under explored. In this paper, we analyse a microservice deployment dataset released by Alibaba to understand its scale, heterogeneity, and dynamicity, and compare our results to previous large-scale deployments to begin to understand their commonalities. We identify tens of thousands of microservices, that support an even broader array of front-end functionality. Moreover, our analysis shows wide-spread long-tailed distributions of characteristics between microservices, such as share of workload and dependencies, highlighting inequality. This diversity is also reflected in call graphs, with front-end service functionalities producing dominant and rarer, non-dominant, call graphs that can involve dissimilar microservice calls. We find that dependencies within the deployment at runtime can be different from the static view of the system, and that the deployment undergoes daily changes. We discuss the implications of our findings for state-of-the-art research in microservice management and research testbed realism.
@@ 1. Introduction
Microservice architectures (MSAs) have become the leading method for building large-scale distributed applications (Livens, 2023). These architectures are built using distinct microservices (MSs) that communicate with each other over well defined APIs. This architectural style facilitates small agile development teams, rapid development of new features, and fine-grained scaling. Yet, these characteristics can also lead to unique challenges associated with MSAs, such as increased operational complexity, MS co-ordination, and scale (Soldani et al., 2018), rendering system maintenance, and integrity and uptime assurance, extremely complicated. However, on the other hand, online application downtime can be associated with extremely high costs, with Fortune 1000 companies reporting losses of up to \$1 million per hour of disruption (Atlassian, 2024; Shultz, 2018).
To address these challenges extensive research has been carried out into novel approaches to detect faults (Nobre, João et al., 2023), pinpoint root causes (Zhang et al., 2025), and provide efficient auto-scaling through workload prediction (Luo et al., 2022). Research has explored data-driven approaches for optimising MSA design and minimising anti-patterns (Gamage and Perera, 2021; Parker et al., 2023). However, such approaches are often motivated and subsequently designed based only on theoretical understandings of MSAs stemming from perceived common design practices and the challenges associated with them. This issue is further exacerbated by the reliance on small-scale open-source testbeds (Seshagiri et al., 2022) due to the scarce access to industrial-scale datasets with ground truth. Whilst such testbeds are designed based on best practices, they do not match the complexity associated with real-world large-scale MSAs (Seshagiri et al., 2022). This puts into question the efficacy of any such derived methodologies.
Previous work has begun to characterise the attributes of large-scale MSAs, such as the properties of MS call graphs (Luo et al., 2021; Wen et al., 2022), runtime performance (Luo et al., 2021; Wen et al., 2022), and the temporal evolution of dependencies at runtime (Winchester et al., 2023). However, these studies only consider characteristics from a narrow viewpoint. To address this limitation, a recent study was carried out on the characteristics of Meta’s MSA (Huye et al., 2023), taking a broader view by investigating the dependency topology, the deployment’s growth over time, and properties of call graphs for several selected front-end functionalities. The study also discusses the implications for research testbeds and how these can be developed to better meet the challenges of real-world deployments. However, the implications of their findings for MS management research are not discussed in depth. Moreover, their investigation omitted several aspects of MSAs that are important for helping the research community understand the challenges associated with these large-scale deployments, such as workload, cohabitation, uncommon call graphs, and runtime topology. Finally, without a substantial comparison to other large-scale deployments, it remains unclear how representative their results are of MSAs as a whole.
The main contributions of this paper are threefold. Firstly, we expand on the analysis carried out in (Huye et al., 2023) by examining three key dimensions of MSAs associated with operational and management challenges from the literature: scale, heterogeneity, and dynamicity (Soldani et al., 2018; Seshagiri et al., 2022; Zhou et al., 2021; Gan et al., 2024, 2021; Lin et al., 2016; Yang et al., 2022; Lin et al., 2018a). Secondly, we offer the first comparative analysis of large-scale MSAs by discussing our findings in light of those from Meta’s deployment. Thirdly, we discuss the implications of our findings for state-of-the-art research in MSA management, design, and testbeds, highlighting where existing work falls short of real-world demands and identifying opportunities for future research.
2. Alibaba’s Microservice Architecture and Dataset
The Alibaba MSA has been previously described in (Luo et al., 2022), for this paper we provide a brief overview. Alibaba have adopted Kubernetes (Kubernetes, 2015) to manage their cloud (Zhang et al., 2020). Stateful MSs, such as databases and in-memory stores, are deployed within a dedicated cluster which is not shared with stateless MSs. Alibaba employ the Application Real-Time Monitoring Service (ARMS) (Alibaba Cloud, [n. d.]; Cai et al., 2019) to collect monitoring information from their deployment. The data collected by ARMS over 14 days of operation of deployed MSs in 2022 has been open-sourced (Alibaba, 2017). The dataset is split into several parts, each covering specific monitoring information; in this paper, we utilise the data associated with MS calls and resource usage, from hereafter referred to as the call data and resource data.
Each row of the call data corresponds to a single call made between an upstream and downstream MS, acting as the consumer and provider MSs, respectively. Each row also contains a number of auxiliary fields. The Trace ID identifies the collection of calls made between MSs that are associated with a single front-end request. The RPC ID is used to reconstruct the call graph structure of a call collection identified by a given Trace ID by establishing the invocation order of calls starting from the entry MS, in a dot-separated format. For example, the two calls made by an entry-level MS A to MSs B and then C will have the RPC IDs ‘0.1’ and ‘0.2’ respectively. If C subsequently made a call to D the RPC ID for that call would be ‘0.2.1’. Finally, the Service ID associates each call group within the data with the invocation of a specific front-end service functionality. These functionalities include searching products, ordering, and delivering [^1] Service IDs are also associated with sporadic non-user functionality such as management and testing, and thus this interpretation of Service IDs is not completely straightforward, as seen in Section 3.2.. However, these are anonymised in the dataset.
Each row of the resource data contains a single value of CPU and memory utilisation from stateless MS instances, sampled every 60 seconds across the 14 days. Each row contains a MS Instance ID and MS ID which correspond to the stateless MS names found in the call data. Each row also contains a Node ID which denotes the specific host in which the MS instance is deployed at the time of the sample.
@@ 3. Scale
MSAs, whilst conceived for individual MS simplicity and modularity, can often evolve into highly complex, sprawling architectures as they organically grow to meet business demands, imparting significant operational challenges for MSA management (Soldani et al., 2018). In this section, we present quantitative insights into how scale manifests at the architectural, service, and MS levels.
@@ 3.1. Microservices and Instances
Scale is measured in tens of thousands of MSs and a million instances: we identified 64,751 unique MSs that were deployed as 1,866,091 instances in 14 days (see Table 1); on average, each MS was deployed as 29 instances.
Most MSs both make and receive calls: we found that 8,591 MSs were responsible for handling entry calls [^2] An entry call is the initial request made from an external client or system that enters the MS deployment. and 25,201 of MSs never made calls across the 14 days, meaning that 30,959 of MSs, both received and made calls. Overall, 61% of leaf MSs were stateful; however, a noteworthy share of leaf MSs were stateless, illustrating that some MSs deliver functionality without any additional database calls.
@@
Alibaba’s deployment is larger in number of unique MSs but Meta’s is larger in number of replications: whilst Meta had fewer unique MSs (18,500 vs 64,800) these MSs were deployed as considerably more instances than at Alibaba (12 million vs 1.8 million). These differences could be a result of the divergent purposes of each deployment, with Meta providing primarily social media services, and Alibaba providing a mix of e-commerce, financial, and digital media services, driving different engineering considerations. Another possibility, which may also be linked to the purpose of each deployment, is that Alibaba opts for much finer-grained MSs, with each MS handling a smaller slice of functionality, whereas Meta consolidates broader functionalities into fewer, more heavily replicated MSs. These results reflect architectural design differences between the two organisations.
Figure 1. Service ID counts vs. the median number of invocations per call fingerprint
@@ 3.2. Services
The number of Service IDs is significantly inflated by sporadic non-user functionality: we identified 166,093,303 unique Service IDs, dwarfing the number of MSs. However, Service IDs can also be associated with non-user functionality. When investigating the number of invocations of each service across the 14 days we found that only 7,352,480 (4.4%) had more than a single invocation, and only 536,958 (0.3%) had more than 100 total invocations. As Alibaba’s online services are generally used by a very large number of users, this suggests that the number of unique services is inflated by non-user facing functionality.
@@
Sporadic Service IDs have call fingerprints that are shared more: for a random sample of 10,000 call fingerprints we grouped Service IDs associated with each call fingerprint and calculated the median invocations. We found that the more a call fingerprint was shared among Service IDs, the fewer invocations each Service ID had within that fingerprint (see Figure 1). Overall, this suggests that a major contributor to Service ID inflation is functionality that appears under different Service IDs across the dataset despite likely representing the same front-end functionality. Therefore, from hereafter we use the Service ID groupings based on these call fingerprints rather than raw Service ID values, which we argue better represent the true repertoire of the front-end functionality provided by the deployment.
@@ 3.3. Workload
Workload significantly fluctuates across each day: the number of calls and invocations fluctuated significantly across the 14 days, with a strong 24-hour periodicity. The average daily trough-to-peak increase was 80% for MS calls and 79% for front-end invocations. The considerable difference between lowest and highest total daily workload highlights the importance of predictive scaling, with inaccurate scaling for peaks and troughs potentially leading to wasted resource usage or impacted user experience respectively.
@@ 3.5. Scale Implications
The number of front-end service functionalities and back-end MSs within an MSA can be extremely large: the number of MSs in the dataset is several orders of magnitude larger than popular testbeds (Gan et al., 2019; FudanSELab, 2018; Platform, 2018; OCP-Power-Demos, 2023). This disparity in scale likely contributes to the limited focus on scalability in current state-of-the-art approaches for MS anomaly detection and localisation. For example, one set of popular approaches construct causal graphs to capture the relationships between MS metrics during a fault (Ma et al., 2020; Chen et al., 2019; Qiu et al., 2020; Wang et al., 2018; Mariani et al., 2018; Wu et al., 2021; Ma et al., 2019, 2021; Lin et al., 2018b; Xin et al., 2023). However, the causal structural modelling methods used by these approaches such as the PC algorithm and LiNGAM, have poor time complexity as the number of variables increases (Spirtes et al., 1993; Colombo and Maathuis, 2014; Shimizu et al., 2006, 2011). Another approach is to train models in a supervised manner, either through the injection of faults to produce a diverse labelled corpus of fault data or through historical root cause information (Zhou et al., 2019; Kohyarnejadfard et al., 2019; Scheinert et al., 2021). However, with such a large number of MSs and diverse set of front-end functionality they support, the feasibility of collecting enough labelled data to adequately cover the large combination of failure conditions is unclear even when leveraging only partial supervised learning (Zhang et al., 2022; Raeiszadeh et al., 2023).
@@
A strategy for reducing challenges associated with characterising the normal behaviour of a MSA is to model the behaviour of individual MSs or MSs associated with individual front-end services (Liu et al., 2020; Yang et al., 2023). However, in this dataset, we identified 8,000 front-end services, therefore training and maintaining a model for each individual front-end service may be costly. Furthermore, we identified a potential 1.2 million distinct front-end functionalities associated with these front-end services. As each front-end functionality is likely associated with its own usage patterns and call graphs, this brings into question whether building a model for each front-end service substantially reduces the problem of scale and diversity. On the other hand, building individual models for all 1.2 million front-end functionalities is unlikely to be feasible outside of very lightweight approaches (Xu et al., 2017).
4. Heterogeneity
MSAs are lauded for their flexibility and scalability (Wolff, 2016; Newman, 2021), achieved through independently developing and deploying loosely coupled MSs. Yet, this independence leads to a diverse and complex ecosystem. In practice, MSA heterogeneity manifests in many dimensions, from the distribution of instances across hosts and the co-location of multiple MSs in shared hardware, to distribution of workloads and the intricate web of dependencies that link them together. Furthermore, MSs can be variably integrated to provide diverse front-end service functionality, generating further avenues for heterogeneity. In this section, we explore how these facets of heterogeneity manifest in this dataset.
4.1. Microservice Instances
Figure 3. Cumulative frequency of the number of instances per MS, with lower and upper quartiles (grey dashed) and median (red dashed).
The number of deployed instances per MS is highly imbalanced: as shown in Figure 3, the distribution of the number of instances per MS is heavy-tailed; 24.8% of MSs were deployed as a single instance, with the median number of instances being 4. On the other hand, the highest number of instances was over 227,000, making up for 12.7% of all deployed MS instances across the 14 days. Furthermore, the top 1% of MSs accounted for around 895,000 instances, making up half of all deployed MS instances. Since we aggregated the number of deployed instances across the full dataset, the number of instances can either mean that a MS is generally deployed with a large number of instances or that a MS is regularly horizontally scaled. However, as explored in Section 5.2, this dataset likely involves core applications with many MSs that do not undergo auto-scaling.
Table 2. Comparison of per‑MS instance counts at Alibaba and Meta

### v3

@@ Abstract.
Microservice management and testbed research often rests on assumptions about deployments that have rarely been validated at production scale. While recent studies have begun to characterise production microservice deployments, they are often limited in breadth, do not compare findings across deployments, and lack consideration of the implications of findings for commonly held assumptions. We analyse a distributed tracing dataset from Alibaba’s production microservice deployment to examine its scale, heterogeneity, and dynamicity. By comparing our findings to prior measurements of Meta’s MSA we illustrate both convergent and divergent properties, clarifying which patterns may generalise. Our study reveals extreme architectural scale, long-tail distributions of workloads and dependencies, highly diverse functionality, substantial call graph variability, and pronounced time-varying behaviour which diverge from assumptions underlying research models and testbeds. We summarise how these observations challenge common assumptions in research on fault management, scaling, and testbed design, and outline recommendations for more realistic future approaches and evaluations.
Cloud Computing, Microservices, Distributed Traces, Management
copyright: noneccs: Networks Cloud computingccs: Networks Logical / virtual topologiesccs: Networks Network manageabilityccs: Computer systems organization Cloud computing
@@ 1. Introduction
Microservice architectures (MSAs) have become the leading method for building large-scale distributed applications (livens_what_2023). These architectures are built using distinct microservices (MSs) that communicate with each other over well defined APIs to provide application logic. This architectural style facilitates small agile development teams, rapid development of new features, and fine-grained scaling. Yet, these characteristics can also lead to unique challenges associated with MSAs, such as increased operational complexity, MS co-ordination, and scale (soldani_pains_2018). This renders system maintenance, integrity, and uptime assurance extremely complicated. However, on the other hand, online application downtime can be associated with extremely high costs, with Fortune 1000 companies reporting losses of up to \$1 million per hour of disruption (atlassian_calculating_nodate; noauthor_cost_nodate).
To address these challenges extensive research has been carried out into novel approaches to detect faults (nobre_anomaly_2023), pinpoint root causes (zhang_failure_nodate), provide efficient auto-scaling through workload prediction (luo_power_2022), and optimise MSA design by minimising anti-patterns (gamage_using_2021; parker_visualizing_2023). However, these approaches are often motivated and subsequently designed based on theoretical assumptions of MSAs stemming from perceived common designed practices and the challenges associated with them. This issue is further exacerbated by the reliance on small-scale open-source testbeds (seshagiri_sok_2022) due to the scarce access to industrial-scale datasets with ground truth. Whilst such testbeds are designed based on best practices, they do not match the complexity associated with real-world large-scale MSAs (seshagiri_sok_2022).
Previous work has begun to characterise the attributes of large-scale MSAs, such as the properties of MS call graphs (luo_characterizing_2021; wen_bytedance_2022), runtime performance (luo_characterizing_2021; wen_bytedance_2022), and the time-varying evolution of dependencies at runtime (winchester_temporal_2023). However, these studies only consider characteristics from a narrow viewpoint. To address this limitation, a recent study was carried out on the characteristics of Meta’s MSA (huye_lifting_2023), taking a broader view of MSA characteristics. However, this analysis still omitted several important characteristics of MSAs and the implications of their findings for the assumptions leveraged by researchers are not discussed in depth. Furthermore, without a substantial comparison to other large-scale deployments, it remains unclear how representative their results are of MSAs as a whole. Despite the growing literature, there still lacks a broad, empirically grounded understanding of how operational complexity manifests at scale, and whether observations at one provider generalise to another.
To address this gap, we make the following three contributions:
  1. (1)
    Comprehensive, multi-dimensional analysis: building on prior work that identifies scale, heterogeneity, and dynamicity as the primary sources of operational complexity in large-scale MSAs (soldani_pains_2018; seshagiri_sok_2022; zhou_fault_2021; gan_sleuth_2024; gan_sage_2021; lin_log_2016; yang_robust_2022; lin_microscope_2018), we carry out a comprehensive, multi-dimensional analysis of Alibaba’s deployment across these three axes, offering a broader view than prior Alibaba and Meta studies.
  2. (2)
    Cross-deployment comparison: we provide the first cross-deployment comparison of two large-scale MSAs, revealing convergent and divergent properties.
  3. (3)
    Operational insights and research implications: we highlight and discuss how real-world MSA operational properties challenge common assumptions in management approaches and testbed design, and we offer recommendations (summarised in Table 4) to guide future work to better meet the challenges of real-world MSA deployments.
All code and the processed data used to generate the figures and statistics in this paper will be publicly released upon acceptance of the paper.
2. Related Work
Quantitative studies of real-world MSAs are relatively few and narrow in scope, though several notable works exist. Luo et al. (luo_characterizing_2021) analysed call graph characteristics from 12 hours of Alibaba traces, reporting depth and complexity statistics and clustering call graphs based on topology. Winchester et al. (winchester_temporal_2023) examined the same data, focussing instead on the time-varying evolution of global dependency structure, revealing repeating system-wide states that could be linked to global system performance. While both studies offered valuable insights, their reliance on a short 12-hour window limits confidence in the completeness of their findings.
Luo et al. (luo_power_2022) later analysed 14 days of Alibaba traces to study time-varying workload in relation to dependency dynamics, topological characteristics, and architectural roles. Despite the larger dataset, this study, similar to previous work, also concentrated on a limited set of characteristics, providing only a partial view of the broader properties of large-scale MSAs.
More recently, Huye et al. (huye_lifting_2023) conducted a broader characterisation of Meta’s MSA, analysing its dependency topology, architectural evolution, and depth and variability of traces for a select few front-end services. Although this work presented a more comprehensive view than previous studies, it does not explore important aspects of MSAs, such as workload, cohabitation, uncommon call graphs, and time-varying topology. Furthermore, this work does not explore the implications of their findings in relation to common assumptions held by MSA management approaches. Moreover, the lack of comparison with other deployments leaves it unclear whether the reported observations reflect common characteristics of large-scale MSAs or instead are specific to Meta’s MSA.
These gaps motivate our comprehensive analysis of Alibaba’s deployment and our cross-deployment comparison with Meta, to exhibit how operational characteristics of real-world MSAs challenge common assumptions in management and testbed design.
Table 1. Example trace from the call data.
Timestamp Trace ID RPC ID Service ID Upstream MS Downstream MS 115352 T_1156086 0.1 S_153587416 MS_58845 MS_71712 115354 T_1156086 0.2 S_153587416 MS_58845 MS_72104 115358 T_1156086 0.2.1 S_153587416 MS_72104 MS_90011
Table 2. Example rows from the resource data.
Timestamp MS ID Instance ID Node ID CPU Util. Mem. Util. 180000 MS_21881 MS_21881_POD_0 NODE_11517 0.21996 0.83300 180060 MS_21881 MS_21881_POD_0 NODE_11517 0.19847 0.81052 180120 MS_21881 MS_21881_POD_1 NODE_11519 0.24216 0.85499
3. Alibaba’s Microservice Architecture and Dataset
The Alibaba MSA has been previously described in (luo_power_2022), for this paper we provide a brief overview. Alibaba have adopted Kubernetes (noauthor_production-grade_nodate) to manage their cloud (zhang_high-density_2020). Stateful MSs, such as databases and in-memory stores, are deployed within a dedicated cluster which is not shared with stateless MSs. Alibaba employ the Application Real-Time Monitoring Service (ARMS) (noauthor_application_nodate; cai_real-time_2019) to collect monitoring information from their deployment. The data collected by ARMS over 14 days of operation of deployed MSs in 2022 has been open-sourced (noauthor_alibabaclusterdata_2025). The dataset is split into several parts, each covering specific monitoring information; in this paper, we utilise the data associated with MS calls and resource usage, from hereafter referred to as the call data and resource data. Certain deployment-specific details that are not publicly available in the description of the dataset and findings from analysis were clarified via private communication with the dataset’s maintainers.
Each row of the call data, shown in Table 1, corresponds to a single call made between an upstream and downstream MS, acting as the consumer and provider MSs, respectively. Each row also contains a number of auxiliary fields. The Trace ID identifies the collection of calls made between MSs that are associated with a single front-end request. The RPC ID is used to reconstruct the call graph structure of a call collection identified by a given Trace ID by establishing the invocation order of calls starting from the entry MS, in a dot-separated format. For example, the two calls made by an entry-level MS A to MSs B and then C will have the RPC IDs ‘0.1’ and ‘0.2’ respectively. If C subsequently made a call to D the RPC ID for that call would be ‘0.2.1’. Finally, the Service ID associates each call group within the data with the invocation of a specific front-end service functionality. These functionalities include searching products, ordering, and delivering [^1] Service IDs are also associated with sporadic non-user functionality such as management and testing, and thus this interpretation of Service IDs is not completely straightforward, as seen in Section 4.2.. However, these are anonymised in the dataset.
Each row of the resource data, shown in Table 2, contains a single value of CPU and memory utilisation from stateless MS instances, sampled every 60 seconds across the 14 days. Each row contains a MS Instance ID and MS ID which correspond to the stateless MS names found in the call data. Each row also contains a Node ID which denotes the specific host in which the MS instance is deployed at the time of the sample.
@@ 4. Scale
MSAs, whilst conceived for individual MS simplicity and modularity, can often evolve into highly complex, sprawling architectures as they organically grow to meet business demands, imparting significant operational challenges for MSA management (soldani_pains_2018). In this section, we present quantitative insights into how scale manifests at the architectural, service, and MS levels.
@@ 4.1. Microservices and Instances
Scale is measured in tens of thousands of MSs and a million instances: we identified 64,760 unique MSs that were deployed as 1,866,091 instances in 14 days (see Table 3); on average, each MS was deployed as 29 instances.
Most MSs both make and receive calls: we found that 8,591 MSs were responsible for handling entry calls [^2] An entry call is the initial request made from an external client or system that enters the MS deployment. and 25,201 of MSs never made calls across the 14 days, meaning that 30,959 of MSs, both received and made calls. Based on the type of RPC IDs of received calls we found that 61% of leaf MSs were stateful. However, a noteworthy share of leaf MSs were stateless, illustrating that some MSs deliver functionality without any additional database calls.
@@
Alibaba’s deployment is larger in number of unique MSs but Meta’s is larger in number of replications: whilst Meta had fewer unique MSs (18,500 vs 64,800) these MSs were deployed as considerably more instances than at Alibaba (12 million vs 1.8 million). This suggests that Meta’s architecture may consolidate broader functionality into fewer MSs that are scaled out more aggressively, whereas Alibaba’s appears to decompose functionality into a larger number of fine-grained MSs. While instance replica count is influenced by many factors such as workload intensity, geo-distribution, and resilience targets, these findings hint at different architectural choices. This divergence likely not only reflects the contrasting service domains of Meta (social media centric) and Alibaba (e-commerce, financial, and digital media) but also differing scaling and operational constraints.
Figure 1. Service ID counts vs. the median number of invocations per call fingerprint.
@@ 4.2. Services
The number of Service IDs is significantly inflated by sporadic non-user functionality: we identified 166,093,303 unique Service IDs, dwarfing the number of MSs. However, Service IDs can also be associated with non-user functionality. When investigating the number of invocations of each service across the 14 days we found that only 7,352,480 (4.4%) had more than a single invocation, and only 536,958 (0.3%) had more than 100 total invocations. However, these values are inflated by non-user facing functionality (see 3). This is reinforced by the fact that Alibaba’s online services are generally used by a very large number of users, thus a large number of services having very few invocations unrealistic.
@@
Sporadic Service IDs have call fingerprints that are shared more: for a random sample of 10,000 call fingerprints we grouped Service IDs associated with each call fingerprint and calculated the median invocations. We found that the more a call fingerprint was shared among Service IDs, the fewer invocations each Service ID had within that fingerprint (see Figure 1). Overall, this suggests that a major contributor to Service ID inflation is functionality that appears under different Service IDs across the dataset despite likely representing the same front-end functionality. Therefore, from hereafter we use the Service ID groupings based on these call fingerprints rather than raw Service ID values, which we argue better represent the true repertoire of the front-end functionality provided by Alibaba’s deployment.
@@ 4.3. Workload
Workload significantly fluctuates across each day: the number of calls and invocations fluctuated significantly across the 14 days, with a strong 24-hour periodicity, see Figure 2. The average daily trough-to-peak increase was 80% for MS calls and 79% for front-end invocations. The considerable difference between lowest and highest total daily workload highlights the importance of predictive scaling, with inaccurate scaling for peaks and troughs potentially leading to wasted resource usage or impacted user experience respectively.
@@ 4.5. Scale Implications
The number of front-end service functionalities and back-end MSs within an MSA can be extremely large: the number of MSs in the dataset (Section 4.1) is several orders of magnitude larger than popular testbeds (gan_open-source_2019; noauthor_fudanselabtrain-ticket_2025; noauthor_googlecloudplatformmicroservices-demo_2025; noauthor_ocp-power-demossock-shop-demo_2025). This disparity in scale likely contributes to management approaches assuming moderate to small deployment sizes. For example, one set of popular approaches construct causal graphs to capture the relationships between MS metrics during a fault (ma_automap_2020; chen_causeinfer_2019; qiu_causality_2020; wang_cloudranger_2018; mariani_localizing_2018; wu_microdiag_2021; ma_ms-rank_2019; ma_servicerank_2022; lin_facgraph_2018; xin_causalrca_2023). However, the causal structural modelling methods used by these approaches have poor time complexity as the number of variables increases (spirtes_causation_1993; colombo_order-independent_2014; shimizu_linear_2006; shimizu_directlingam_2011). The implicit reliance on this assumption challenges the suitability of many prior approaches to real-world large-scale systems.
This disparity in deployment size also leads to another common assumption that an adequately representative labelled corpus can be collected through fault injection or historical root cause information (zhou_latent_2019; kohyarnejadfard_system_2019; scheinert_learning_2021). However, with such a large number of MSs and diverse set of front-end functionality they support, the feasibility of collecting enough labelled data to adequately cover the large combination of failure conditions is unclear even when leveraging only partial supervised learning (zhang_putracead_2022; raeiszadeh_real-time_2023).
@@
A strategy for reducing challenges associated with characterising the normal behaviour of a MSA is to model the behaviour of individual MSs or MSs associated with individual front-end services based on the assumption that per-front-end service modelling is more tractable (liu_unsupervised_2020; yang_transparently_2021). However, in this dataset, we identified 8,000 front-end services (Section 4.2), therefore training and maintaining a model for each individual front-end service may be costly. Furthermore, we identified a potential 1.2 million distinct front-end functionalities associated with these front-end services. As each front-end functionality is likely associated with its own usage patterns and call graphs, this challenges whether building a model for each front-end service substantially reduces the problem of scale and diversity. On the other hand, building individual models for all 1.2 million front-end functionalities is unlikely to be feasible outside of very lightweight approaches (xu_lightweight_2017).
The workload experienced by the system fluctuates significantly across the day: widely used testbeds generally use stationary workload generators (tene_giltenewrk2_2024) to simulate user behaviour. This means that generated datasets only represent system behaviour under stationary conditions (pham_baro_2024; li_dejavu_2022; lee_eadro_2023; zhang_deeptralog_2022). Subsequently, management approaches being developed and benchmarked using these deployments implicitly assume that workload, metrics, and call graphs are stationary. Our results (Section 4.3) reveal strongly fluctuating user demand, indicating that methods trained and evaluated under stationary conditions may not generalise well to real-world variable workloads. This is particularly true for approaches that leverage change point detection for anomaly alerting (pham_baro_2024).
(a) CDF of instances per microservice (Alibaba).
@@
(b) Per-MS instance counts comparison.
Figure 3. Alibaba’s instance distribution and comparison with Meta: (a) CDF of instances per microservice with with lower and upper quartiles (grey dashed) and median (red dashed); (b) summary statistics highlighting the contrasting instance counts.
5. Heterogeneity
MSAs are lauded for their flexibility and scalability (wolff_microservices_2016; newman_building_2021), achieved through independently developing and deploying loosely coupled MSs. Yet, this independence leads to a diverse and complex ecosystem. In practice, MSA heterogeneity manifests in many dimensions, from the distribution of instances across hosts and the co-location of multiple MSs in shared hardware, to distribution of workloads and the intricate web of dependencies that link them together. Furthermore, MSs can be variably integrated to provide diverse front-end service functionality, generating further avenues for heterogeneity. In this section, we explore how these facets of heterogeneity manifest in this dataset.
The number of deployed instances per MS is highly imbalanced: as shown in Figure 3(a), the distribution of the number of instances per MS is heavy-tailed. 24.8% of MSs were deployed as a single instance, with the median number of instances being 4. On the other hand, the highest number of instances was over 227,000, making up for 12.7% of all deployed MS instances across the 14 days. Furthermore, the top 1% of MSs accounted for around 895,000 instances, making up half of all deployed MS instances. Since we aggregated the number of deployed instances across the full dataset, the number of instances can either mean that a MS is generally deployed with a large number of instances or that a MS is regularly horizontally scaled. However, this dataset includes MSs from core applications that undergo much less auto-scaling (see 3), this can be observed in our findings in Section 6.2.
MSs at Alibaba and Meta demonstrate similar distributions of, and maximum, instance counts: the deployed instances per MS at both MSAa demonstrate a heavy-tail (see Figure 3(b)), however, the median, average and 99-th percentile values for Meta are an order of magnitude higher. Given that the number of instances deployed at Meta is also an order of magnitude higher (see Section 4.1), this means that the central tendencies for both distributions when normalising by the total number of instances are similar, and thus the typical MS within both deployments represents a similar fraction of the overall deployed instances. Notably, we found that the maximum number of instances for a single MS were remarkably similar (227,000 vs 557,000) despite this order of magnitude difference in the total number of instances. This indicates that the heavy tail is more pronounced in this dataset, with the MS with the most instances accounting for 12.4% of the total number of instances compared to Meta’s 4.6%. The convergence of maximum values despite the difference in the total number of instances might suggest that there are common operational ceilings when it comes to scaling. This ceiling could be a result of coordination overhead, such as managing load balancing between a larger number of instances, or non-linear performance gains due to resource contention for shared resources such as databases, caches, and network bandwidth.